I need to find the indicated outputs for
f(x)= 3x^2-2x
f (0) f (-1) f (2)
Thank you

Answer by Alan3354(69443)

I need to find the indicated outputs for
f(x)= 3x^2-2x
f (0) f (-1) f (2)
---------------
f(x) = 3x^2 - 2x
For f(-1) sub -1 for x
f(-1) = 3(-1)^2 - 2(-1) = 3 + 2 = 5
---------------
Do the others the same way.